OpenStudy (anonymous):

What is the transformation that occurs to the equation y=2^x if the equation changes to y= 2^x-3+4.

OpenStudy (anteater):

Is your equation y = 2^x - 3 + 4 or \[2^{x-3} + 4\] ?

OpenStudy (anteater):

I am presuming it is the second one, which would shift the graph right 3 units and up 4.

OpenStudy (mathmale):

anteater is right in asking exactly what you meant by y=2^x-3+4. Without parentheses, you have in effect written y=2^x+7. I'm sure that's not the function you found in your homework assignment. It is absolutely essential that you enclose that (x-3) in parentheses. If you do that, you end up with \[y=2^{(x-3)}+4.\]

OpenStudy (mathmale):

this tells the reader that the exponent here is (x-3), and not just x.

OpenStudy (mathmale):

Now, if you begin with y = 2^x, and graph it, you'll find that the outcome is an exponential function whose y-intercept is 1. If you (1) translate this graph 3 units to the right, (2) as well as 4 units up, you have in fact gone from y = 2^x to y = 2^(x-3) + 4.